### v1

# A Systems Thinking Approach to Algorithmic Fairness

Abstract.
Systems thinking provides us with a way to model the algorithmic fairness problem by allowing us to encode prior knowledge and assumptions about where we believe bias might exist in the data generating process. We can then model this using a series of causal graphs, enabling us to link AI/ML systems to politics and the law. By treating the fairness problem as a complex system, we can combine techniques from machine learning, causal inference, and system dynamics. Each of these analytical techniques is designed to capture different emergent aspects of fairness, allowing us to develop a deeper and more holistic view of the problem. This can help policymakers on both sides of the political aisle to understand the complex trade-offs that exist from different types of fairness policies, providing a blueprint for designing AI policy that is aligned to their political agendas.
systems thinking, machine learning, causal inference, system dynamics, emergence, bias, fairness, discrimination, disparate treatment, disparate impact
copyright: acmlicensedjournalyear: 2025doi: XXXXXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; 2025; Athens, Greeceisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Causal reasoning and diagnosticsccs: Computing methodologies Probabilistic reasoningccs: Computing methodologies Reasoning about belief and knowledgeccs: Applied computing Lawccs: Applied computing Sociologyccs: Applied computing Economicsccs: Applied computing Psychologyccs: Social and professional topics Governmental regulations
[…]
Figure 1. Three different representations of algorithmic fairness
[…]
A block diagram, causal Bayesian network, and causal loop diagram.
1. Introduction
For several years, AI researchers have struggled with how to build AI/ML systems that are fair and aligned to society’s values. Purely technical approaches that rely on statistical correlation alone have fallen far short of what is needed to solve this difficult problem. Systems thinking offers a novel approach to the algorithmic fairness problem by allowing us to unify complex ideas from the social sciences and computer science into a comprehensive and integrated theory. By leveraging prior knowledge and assumptions about where we believe bias might exist in the data generating process, we can help policymakers on both sides of the aisle to design AI policy that is aligned to their political agendas.
To address the algorithmic fairness problem, we need to take a sociotechnical approach. We will begin by analyzing the social aspects of this problem using first principles. We will explore the nature of free will to develop an understanding of moral responsibility. We will review the structure versus agency debate in sociology and attribution theory in psychology to address the controversial question: Why do we believe disparities exist across protected groups like race or gender? This will give us insights into the sources of bias in the data generating process, showing us what different interventions are needed to make a decision fair. We can then use this knowledge to build a better understanding of the political debate around fairness, enabling us to develop a deeper and more holistic way to analyze this complex and difficult problem.
We will then turn our attention to the technical aspects of the fairness problem, which forms the bulk of this paper. We will combine techniques from machine learning, causal inference, and system dynamics to develop different causal models of the fairness problem. These three analytical techniques model different facets of the fairness problem and are complementary to each other. Machine learning is necessary to build AI systems that can make fair predictions. Causal inference allows us to model bias from the data generating process in an acyclic manner, allowing us to visualize different forms of discrimination and what types of interventions are needed to make fair decisions. System dynamics allows us to model bias from the data generating process in a cyclic manner, allowing us to capture counterintuitive behavior in a complex system to explain why disparities might exist across protected groups. The integration of these three previously independent analytical techniques could also improve our understanding of causality, providing a new causal hierarchy for modeling fairness.
Finally, we will bring together the social and technical aspects of the fairness problem. We will build a system map to show how the social sciences and computer science are related to each other. We will show the dualistic nature of the fairness problem across the social sciences and how this relates to Pearl’s two fundamental laws of causal inference. Hopefully, this will help us to finally bridge the sociotechnical fairness gap.
2. Related work
The field of algorithmic fairness is a relatively new field that has drawn significant attention since the release of a ProPublica article that claimed racial bias in COMPAS, a software tool that is widely used in the US court system to predict recidivism(Angwin et al., 2016). Since then, there have been multiple attempts to align the legal definitions of discrimination (e.g. disparate treatment and disparate impact in the US and direct discrimination and indirect discrimination in the EU) with their equivalent mathematical definitions. But these approaches, which have oftentimes relied on statistical correlation, have not been successful. Narayanan for example identified nearly two dozen fairness metrics, none of which were aligned well to the legal definitions of discrimination(Narayanan, 2018).
Pearl has argued for a few decades that the discrimination problem is fundamentally a causality problem(Pearl, 2001). He has argued that fairness can be modeled using causal Bayesian networks (CBNs) using a series of direct and indirect effects. This work was later expanded upon by Plecko and Bareinboim, who proposed the idea of a standard fairness model(Plecko and Bareinboim, 2022). They argued that there exists a causal template through which entire classes of fairness models can be derived using a small number of modeling assumptions. Barocas et al. devoted a chapter in their textbook to discuss how causality might be used to model fairness(Barocas et al., 2023). Kusner et al. argued for the use of counterfactuals in modeling fairness(Kusner et al., 2017). Chiappa and Isaac showed that certain paths within a CBN could be labeled as fair or unfair(Chiappa and Isaac, 2019). Another attempt of building causal fairness models was made by Nilforoshan et al., but their approach had encountered limitations(Nilforoshan et al., 2022).
However, their causal approaches were not really sociotechnical and didn’t take full account of the social and political aspects of the fairness problem. Society has different ideological beliefs about what is fair and unfair based on different sets of assumptions about where bias is in the data generating process. Our paper addresses this issue by encoding society’s different mental models into causal models using first principles. By taking into account a diversity of perspectives, we can understand the politics behind the fairness problem. This understanding of the politics is critical for linking machine learning to the law.
Another key difference is that their approaches relid solely on causal inference, which has limitations in its ability to model the data generating process. Our approach leverages systems thinking and specifically system dynamics to gain a fuller understanding of the data generating process through the modeling feedback loops. Furthermore, we link supervised machine learning to causal inference by recognizing that a CBN can be used as a higher level abstraction of a black box ML model. By bringing machine learning, causal inference, and systems thinking together, we can model the different aspects of the algorithmic fairness problem.
3. Towards a unified theory of fairness
To build fair AI/ML systems, we begin from first principles with the strong assumption that all humans have free will(Carus, 1910). This makes us fundamentally different from machines, which are programmed by humans and therefore have no free will of their own (at this time).
Free will forms the basis of moral responsibility(Talbert, 2024). If we humans have free will, then we have the agency to make our own decisions. These decisions will lead to certain outcomes, and we may be held morally responsible and accountable for those outcomes. Since machines have no free will, they cannot be held morally responsible for any decisions that they make(Grodzinsky et al., 2008). However, the humans who programmed the machines can still be held accountable(Kuflik, 1999; Hew, 2014).
Most of the time, we humans make decisions that only affect our own selves. We can be held personally responsible for the decisions that we make for ourselves. But sometimes, we humans make decisions that affect other people. This is especially true for humans who are in positions of great power. We may be held socially responsible for the decisions that we make for others.
Sometimes the decisions that we make will lead to good outcomes for other people. In these cases, they might assign us praise and reward. On the other hand, sometimes the decisions that we make will lead to bad outcomes for other people. In these cases, they might assign us blame and punishment. This concept is known as moral desert(Kagan, 2012). We humans have a natural tendency to accept praise and reward, and to reject blame and punishment.
But sociologists would note that we are never fully free to make our own decisions. This is because social structures can influence our agency(Barker, 2005). Social structures define our relationships with the various institutions of society. These could include our economic systems, legal frameworks, cultural norms, and political institutions. Agency defines our ability to act independently and to freely make our own decisions. Social structures place a limit on the the choices and opportunities that are available to us humans. Although we may want to live in a society where everyone has equal access to opportunities(Elford, 2023), the reality is that some humans have access to more opportunities than others. This debate between structure versus agency is central to our understanding of human behavior.
To better understand our behavior, we turn to a concept in psychology called the locus of control. This was coined by Rotter and refers to the degree of belief that people have control over the outcomes of events in their own lives(Rotter, 1966). Humans with an internal locus of control believe that they are in control of their own lives, while those with an external locus of control believe that their lives are controlled by outside factors.
How structure impacts agency leads us to another concept in psychology called learned helplessness. Pioneered by Seligman, the idea is that people who have been repeatedly exposed to challenges may believe that they have no control over improving their situation(Seligman, 1972). As a result, they give up trying and have learned to accept their fate, resulting in an external locus of control. For example, disadvantaged groups who grow up in poor neighborhoods are repeatedly exposed to challenges that are outside of their control, reducing their sense of agency over the long term. This could then lead to a victimhood mindset(Kaufman, 2020), which is a tendency for humans to blame others for their misfortunes even if the fault is sometimes their own.
A key concept for understanding fairness in psychology is attribution theory, which is used by humans to interpret and assign causes for events. First proposed by Heider(Heider, 1958) and expanded by Kelley(Kelley, 1967) and Weiner(Weiner, 1986), attribution theory examines whether attributions should be assigned to internal (dispositional) attributions or external (situational) attributions. The internal attributions may be due to behavior or personal traits, while the external attributions may be due to situational or environmental factors. This is essential for determining whether we believe an outcome is biased and therefore unfair.
For example, let’s say that a student did poorly on an exam. An internal attribution might argue that the student was lazy and should have spent more time studying. In this case, the student is to blame for the poor results and thus the test was unbiased and fair. An external attribution might argue that the test was too hard or that the material was not taught well. In this case, the professor is to blame for the poor results and thus the test was biased and unfair. To determine whether an outcome is biased and unfair, all we need to know is whether we would attribute a bad outcome as being internal or external. We can think of algorithmic fairness as a type of credit assignment problem. But since we usually focus on bad outcomes, this becomes more of a blame assignment problem.
The fact that fairness is ultimately based off of normative beliefs instead of positive truths can lead to all sorts of cognitive biases. One source of bias is the fundamental attribution error, where we might wrongly underemphasize external attributions and overemphasize internal attributions(Ross, 1977). For example, if somebody cuts you off while you’re driving, you might blame them for being reckless when in fact they might have been in a rush to bring a dying patient to the hospital. Another type of bias is self-serving bias, where we might try to protect our own self-esteem by assigning internal attributions to our own success and external attributions to our own failures(Myers, 2017). For example, we might attribute winning a game to our own talent and losing a game to an opponent’s cheating.
Therefore, we encounter a causality dilemma when we examine the outcomes of individuals or groups. When an individual receives a bad outcome, is it the fault of the individual’s own actions (an internal attribution) or an unjust situation (an external attribution)? This can become a particularly heated and controversial discussion when we examine outcomes at the aggregate group level. Why have disadvantaged groups received worse outcomes than advantaged groups? Are they victims of an unjust system that has denied them opportunities (an external attribution), or have they adopted a victimhood mindset where they have been conditioned to be helpless (an internal attribution)?
This difference in attribution would lead to a difference in the assignment of moral responsibility. Those who attribute disparities as being external might be more likely to frame this as a social responsibility problem. That is, they would argue that unjust social structures are responsible for creating the disparities across groups. Therefore, disadvantaged groups cannot be held responsible for their outcomes because they did not have access to opportunities. In contrast, those who attribute disparities as being internal might be more likely to frame this as a personal responsibility problem. That is, they would argue that poor individual choices are responsible for creating the disparities across groups. While social structures can limit opportunities, people should still be held personally responsible for making good choices given whatever opportunities they do have access to.
These differences could also affect our perception of inequality. Those who attribute disparities as being external might view inequality as more of a “bug” of the system. It may reflect historically unjust policies that systematically denied resources from the groups who are most in need. They may hold more negative views of capitalism, which naturally creates and reinforces inequalities across groups(Piketty and Goldhammer, 2014). On the other hand, those who attribute disparities as being internal might view inequality as more of a “feature” of the system. It may motivate individuals to make better choices for themselves through the unequal incentives provided by the invisible hand of capitalism(Smith, 1759, 1776). They may hold more positive views of capitalism as an engine for prosperity by providing the motivation for people to work hard.
The two sides might prescribe to different and opposing philosophical schools of thought. Those who attribute disparities as being external might be more likely to prescribe to Rawl’s theory of justice(Rawls, 1999), which argues that all humans before they have even been born start from an original position of equality. From this position, our principles of justice should be chosen behind a veil of ignorance, meaning that we should ignore our own personal circumstances. Since we wouldn’t know in advance whether we would be born into the least advantaged class, we would want society to institute an economic system that benefits the least advantaged. This is known as Rawl’s difference principle. However, those who attribute disparities as being internal might be more likely to prescribe to Rand’s theory of objectivism(Peikoff, 1993). She argues that reality is objective – it is what it is and not what we want it to be. She also argues that reality is independent of one’s consciousness, so we should not imagine that we all start from Rawl’s original position of equality. Therefore, we should accept reality for what it is and then make the most out of it given our own personal circumstances.
When it comes to racial disparities, we would also see a divide in philosophical perspectives. Those who attribute racial disparities as being external might be more likely to prescribe to Bell’s critical race theory, which argues that disparities are primarily driven due to unjust social structures(Delgado and Stefancic, 1998). Historical policies like slavery, Jim Crow segregation, and redlining created social inequalities that continue to persist today. They might also be proponents of Crenshaw’s intersectionality framework, which examines how overlapping identifies like race and gender could compound discrimination or privilege(Crenshaw, 1991, 1989). They might argue in favor of racially affirmative policies to correct for those past injustices. But those who attribute racial disparities as being internal might be more likely to prescribe to Sowell’s theories, which argue that individual agency contributes to the disparities(Sowell, 2019). They might argue that the disparities can be due to differences in culture, beliefs, and practices. They might also argue that racially affirmative policies could cause more harm than good, and therefore argue in favor of racially blind policies that reward people based solely on their merit and qualifications(Sowell, 2004).
This distinction between external and internal attributions can help explain the differences in preference between group fairness and individual fairness(Mitchell et al., 2021; Castelnovo et al., 2022; Verma and Rubin, 2018; Mehrabi et al., 2021; Binns, 2019). Those who attribute disparities as being external may prefer group fairness, meaning that different groups (regardless of differences in average qualifications) should receive similar outcomes on average. After all, there should be no differences between groups because they cannot be blamed for any disparities. Those who attribute disparities as being internal may prefer individual fairness, meaning that similarly qualified individuals (regardless of group membership) should receive similar treatment. That is, individuals who want better outcomes should simply work harder so that they can prove themselves through their own merit and qualifications.
Finally, the divide between internal versus external attribution forms the basis of the political and legal debate over how to close the disparities between protected groups(Sowell, 2019; Donald, 2023; Mills, 1997; Kendi, 2017; Loury, 2002). Those on the left-side of the political spectrum (e.g. liberals and progressives) are more likely to attribute disparities as being external. This would make them believe that the system and the data that it produces is fundamentally biased and unjust. They might argue for large government interventions like affirmative action or DEI to correct for social injustices. They might focus more on preventing disparate impact (also known as indirect discrimination), a form of illegal discrimination where a neutral policy may disproportionately impact one group more than another. Meanwhile, those on the right-side of the political spectrum (e.g. conservatives and libertarians) are more likely to attribute disparities as being internal. This would make them believe that the system and the data that it produces is fundamentally unbiased and just. They might argue for limited government intervention like fairness through unawareness. They might focus more on preventing disparate treatment (also known as direct discrimination), a form of illegal discrimination where individuals are being treated differently based on their membership in a protected group.
This gives us a basic philosophical framework for understanding society’s different ideological viewpoints about whether a system is biased and unfair. We will now encode these ideas for modeling our different beliefs about bias from the data generating process into several causal graphs. By understanding society’s diverse and oftentimes contradictory beliefs about the origins of bias, we can use these graphs to model different forms of discrimination. This will allow us to understand the different types of interventions people might want to employ to ensure that a decision is fair.
4. Modeling fairness as a complex system
The algorithmic fairness problem can be modeled as a complex system by merging techniques from machine learning, causal inference, and system dynamics. Each of these techniques capture different aspects of the fairness and can be seen as operating at different layers within a larger causal hierarchy. The first layer is machine learning, where we want models to make fair predictions and decisions that are compliant with society’s laws and regulations. The next layer is causal inference, where we can model different forms of bias and discrimination. This helps us to identify what interventions are necessary to make a fair decision. The final layer is system dynamics, where we can model the dynamics of the fairness problem to identify counterintuitive behavior within the system. This allows us to build mental models for how different fairness interventions can affect disparities across protected groups.
Figure 1 provides a visual representation of these three different but complementary analytical techniques. Notice that in each representation, we try to use the same variables to demonstrate that they are different ways of looking at the same problem. Examples of what each of the variables in these figures might represent are listed for several high-stakes applications in Table 1.
Table 1. Examples of high stakes applications for algorithmic fairness
Application Protected Attribute $A$ Mediator $W$ Data $X$ Decision $D$ Outcome $Y$
Credit Race, gender Creditworthiness Income, credit history Deny loan? Loan default?
Insurance Race, gender Insurability Demographics, claims history Deny policy? Claim filing?
Housing Race, gender Home value Location, size, condition Home valuation? Final sales price?
Employment Race, gender Qualifications Experience, education Screen resume? Turnover?
Education Race, gender Merit Grades, test scores Reject applicant? Graduation?
Research Race, gender Merit Quality, clarity, originality Reject paper? No. of citations?
[…]
A table defining examples of variables for several high stakes applications.
We begin with Figure 1a, which is a block diagram representation of supervised machine learning. In the center of this block diagram is a black box machine learning model that receives data $X$ as input and returns decision $D$ as output. The decision $D$ leads to some outcome $Y$, which becomes future data $X$ through a feedback loop. A key question is whether a protected attribute $A$ should or should not be fed into the black box model and if so, then how.
Next up we have Figure 1b, which is a causal Bayesian network representation (CBN) of causal inference. CBNs were pioneered by Pearl(Pearl, 2000; Pearl and Mackenzie, 2018; Pearl et al., 2016) and allow us to symbolically encode knowledge and assumptions about causal effects between variables, which are shown in this graph as nodes. The first node of the CBN represents a protected attribute $A$, which has a direct effect on a mediator $W$ ($A\rightarrow W$). The mediator $W$ has a direct effect on an outcome $Y$ ($W\rightarrow Y$). We assume that the mediator $W$ completely explains away the causal effect between the protected attribute $A$ and the outcome $Y$ ($Y\perp\!\!\!\perp A|W$). However, the mediator $W$ might be considered biased by some people if they attribute the disparities across groups as being external. The mediator $W$ is filled white because it is latent, but it can still be partially measured as data $X$ ($W\rightarrow X$). Data $X$ is then fed into a machine learning model that is embedded inside the node for decision $D$ ($X\rightarrow D$). Finally, the decision $D$ also has a direct effect on an outcome $Y$ ($D\rightarrow Y$). CBNs are represented by directed acyclic graphs (DAGs), which means that we cannot use them to model feedback loops.
Finally we have Figure 1c, which is a causal loop diagram (CLD) representation of system dynamics. CLDs were pioneered by Forrester (Forrester, 1971) and are similar to CBNs in that they allow us to symbolically encode knowledge and assumptions about cause and effect relationships between variables(Sterman, 2000). But unlike CBNs, CLDs allow us to model feedback loops. This is useful for modeling why we might believe that disparities exist between protected groups. In the center of the CLD is a node for decision $D$, which like before contains our machine learning model. This CLD is two sided, with the left side representing group A and the right side representing group B. Data from group A $X_{A}$ and group B $X_{B}$ is fed into the ML model, which renders a decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). In this scenario, a decision $D$ that leads to a more favorable outcome for group A $Y_{A}$ might also lead to a less favorable outcome for group B $Y_{B}$ ($Y_{A}\leftarrow D\rightarrow Y_{B}$). This could happen on an absolute basis if both groups are competing for the same resources (e.g. a seat at an elite university). This could also happen on a relative basis if one group receives proportionately more favorable decisions than another (e.g. one group receives higher loan approval rates than another). Note that there is a time delay (denoted by ——) between decision $D$ and outcome $Y$. The outcome $Y$ is recorded as future data $X$ ($Y_{A}\rightarrow X_{A}$ and $Y_{B}\rightarrow X_{B}$), which in turn drives a future decision $D$ ($X_{A}\rightarrow D\leftarrow X_{B}$). We represent this using two reinforcing feedback loops (R1 and R2), which we will show later results in a recurring pattern of behavior called the “success to the successful” archetype.
Let’s say that group A has historically been more successful than group B, which was due to some unjust practice from the past. Then when a model is making a decision $D$ about how to allocate resources to individuals based on past data $X_{A}$ and $X_{B}$ ($X_{A}\rightarrow D\leftarrow X_{B}$), it may allocate more resources to individuals from group A since it would be using that group’s past success to predict future success. This could breed resentment among members of group B, who might demand the use of their protected attribute $A_{B}$ to intervene on a decision $D$ ($A_{B}\rightarrow D$). Therefore, rather than the ML model making a decision $D$ on data $X$ alone, it might use the protected attribute $A_{B}$ to affirmatively favor group B. However, this affirmative action policy might have an unintended consequence that could harm group B’s future outcomes $Y_{B}$ ($A\rightarrow Y_{B}$), offsetting the benefits of an affirmative action policy. We represent this for both sides using a series of balancing feedback loops (B1 and B2) and reinforcing feedback loops (R3 and R4), which we will show later results in a recurring pattern of behavior called the “shifting the burden / addiction” archetype.
These causal graphs form the basis for modeling the complex political and legal debate over how society can build fair machine learning models. They allow us to model the social structures that influence human behavior over time. In the next several sections, we will dive deeper into how to model bias and discrimination using these causal graphs. They will allow us to understand the complex tradeoffs that exist from different fairness policies, providing a set of mathematical rules for policymakers on how to design AI policies to achieve their political goals.
5. Modeling fairness as a causal Bayesian network
Causal Bayesian networks (CBNs) can be used as a higher level representation of any supervised machine learning model. It is at this layer of abstraction that we can begin modeling the different biases that society might have. By understanding these various sources of bias, we can model the different ways that a model can discriminate. Finally, we can use these insights to help policymakers on both sides of the aisle to design AI policies for machine learning models that they would deem to be fair.
Figure 2 shows a series of CBNs modeling different forms of bias, fairness, and discrimination. Each CBN captures a different way of framing the fairness problem based on an individual’s prior beliefs about where bias might exist in a model. On the left, we have four different forms of fairness, which we will later show can be mapped to different points on the political spectrum. On the right, we have four different forms of discrimination, which can be mapped directly to antidiscrimination law.
(a) Fairness through supremacism $A$$W$$X$$D$$Y$ $A$$W$$X$$D$$Y$ (b) Overt discrimination
(c) Fairness through unawareness $A$$W$$X$$D$$Y$ $A$$W$$X$$D$$Y$ (d) Covert discrimination
(e) Fairness through affirmative action $A$$W$$X$$D$$Y$ $A$$W$$X$$D$$Y$ (f) Reverse discrimination
(g) Fairness through lottery $A$$W$$X$$D$$Y$ $A$$W$$X$$D$$Y$ (h) “No” discrimination
Figure 2. Causal Bayesian network representation of algorithmic fairness
[…]
A series of causal Bayesian networks modeling different forms of bias, fairness, and discrimination.
The first form of fairness that we will look at is fairness through unawareness (Figure 2c). This is the simplest form of fairness. We begin with the belief that the data is fundamentally unbiased (internal attribution), which is why all of the nodes in the graph are colored black. In this form of fairness, there is no edge between the protected attribute $A$ and the decision $D$. Therefore, the model is blind to an individual’s membership to a protected group $A$ and therefore cannot treat any individual differently on that basis. We would say that the protected attribute $A$ has no direct effect on the decision $D$, making the decision fair.
The advantage of fairness through unawareness is that it avoids overt discrimination (Figure 2b). If a model uses knowledge of an individual being a member of a disadvantaged group against them in its decision $D$, then it would clearly cause disparate treatment or direct discrimination. We show this as a form of bias where the edge from the protected attribute $A$ to decision $D$ is colored red ($A\rightarrow D$). This leads to a biased decision $D$, which in turn causes a biased outcome $Y$ ($D\rightarrow Y$). Note that in these CBNs, bias is represented by colored nodes and flows across the entire graph. We would say that the protected attribute $A$ has a negative direct effect on the decision $D$ ($A\rightarrow D$), leading to harmful discrimination.
But if we instead believe that the data is fundamentally biased (external attribution), then fairness through unawareness could lead to covert discrimination (Figure 2d). In this scenario, there is bias between the protected attribute $A$ and the mediator $W$ ($A\rightarrow W$), which flows across the entire graph and causes all of the nodes in the graph to be colored red. This bias can be attributed to structural or systemic forms of discrimination that unfairly disadvantages one group over another. For example, if a disadvantaged group was overtly discriminated against in the past, then the present data $X$ may still reflect that negative bias. By removing the edge from the protected attribute $A$ to the decision $D$, it could reinforce one group’s prior success over another. Covert discrimination could form the basis of a disparate impact or indirect discrimination claim. We would say that the protected attribute $A$ has a negative indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$), leading to harmful discrimination.
To correct for covert discrimination, one might choose fairness through affirmative action (Figure 2e). In this form of fairness, the protected attribute $A$ is used in the decision $D$ ($A\rightarrow D$) to correct for bias from the data $X$. The intervention by the protected attribute $A$ is colored blue, which offsets the bias from the data $X$ that has been colored red. As a result, the final decision $D$ is colored black to indicate that its bias has been corrected. Interestingly, the outcome $Y$ is still colored red as the protected attribute $A$ does not have a direct effect on the outcome $Y$. The protected attribute $A$ instead has an indirect effect on the outcome $Y$ through the decision $D$, but this tends to be a weaker effect. This intervention does not directly fix the underlying biases in the mediator $W$, which has a strong direct effect on the outcome $Y$ ($W\rightarrow Y$). We would say that the protected attribute $A$ has a negative indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$) and a positive direct effect on the decision $D$ ($A\rightarrow D$), thus canceling out the biases and making the decision fair.
But if we disagree with the belief that the data is fundamentally biased (internal attribution), then fairness through affirmative action could be perceived as a form of reverse discrimination (Figure 2f). The intervention of the protected attribute $A$ on the decision $D$ ($A\rightarrow D$) could lead to individuals receiving a positive decision $D$ but a negative outcome $Y$. For example, a machine learning model may predict that an individual from a historically disadvantaged group may end up with a negative outcome $Y$. However, this would lead to a greater disparity in decisions across protected groups. So the model may instead use knowledge of that individual’s membership in a protected group $A$ to render a more positive decision $D$. But this positive decision $D$ could lead to a worse outcome $Y$ for the individual than if that individual had just received a negative decision $D$ in the first place. An example might be approving a minority borrower for a loan even if that borrower is not creditworthy, which could result in a loan default. We would say that the protected attribute $A$ has a positive direct effect on the decision $D$ ($A\rightarrow D$), leading to harmful discrimination.
We now move on to the more extreme forms of fairness. Reverse discrimination could lead to fairness through supremacism (Figure 2a). An advantaged group may use this approach to actively suppress a disadvantaged group (e.g. white supremacism). This could happen if an advantaged group believes that a disadvantaged group has been unfairly receiving preferential treatment. For example, if a disadvantaged group has benefited from reverse discrimination in the past, then the present data $X$ may still reflect that positive bias and is colored blue. The protected attribute $A$ is then used to correct for that positive bias in the decision $D$ ($A\rightarrow D$) and is colored red. This could explain what happened when hiring managers racially discriminated against job applicants with Black sounding names compared to those with White sounding names even through the resumes were the same, as shown by Bertrand and Mullainathan(Bertrand and Mullainathan, 2003). We would say that the protected attribute $A$ has a positive indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\rightarrow D$) and a negative direct effect on the decision $D$ ($A\rightarrow D$), thus canceling out the biases and making the decision fair.
Another extreme form of fairness is fairness through lottery (Figure 2g). If we believe that the data is fundamentally biased, an alternative approach to removing bias in the decision $D$ is to remove the edge from data $X$ to decision $D$. The advantage of this approach to fairness is that it provides stronger guarantees of equalizing outcomes across groups. This is particularly effective for dealing with intersectionality issues, ensuring that every potential intersection of a protected class (e.g. race, gender, and religion) receives the same decision rates. Individuals cannot change the decision by altering their own data $X$. The decision $D$ could be assigned randomly or there could be a fixed rule where everyone receives a positive decision $D$ or a negative decision $D$. We would say that the protected attribute $A$ has no direct effect on the decision $D$ and its indirect effects have been canceled out, making the decision fair.
But fairness through lottery could be interpreted as “no” discrimination (Figure 2h). Since data $X$ has no effect on the decision $D$, then the machine learning model can literally not discriminate any differences between individuals. In doing so, it cannot cause illegal discrimination either. But this could still be interpreted as a decision that is positively biased in favor of a disadvantaged group. We would say that the protected attribute $A$ has a positive indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\not\rightarrow D$), leading to harmful discrimination.
CBNs are a very effective way to model the normative effects of fairness and discrimination. They allow us to encode our beliefs about how we might believe that AI systems ought to behave.
Table 2 lists the normative effects of each form of fairness as a function of how the protected attribute $A=a$ affects a decision $D=d$. The total effect is the sum of the direct and indirect effects: $TE_{a}(d)=DE_{a}(d)+IE_{a}(d)$. By definition, the total effect must be equal to zero for it to be considered fair: $TE_{a}(d)=0$.
Table 2. Normative effects of fairness
Effect Lottery Affirmative action Unaware- ness Suprema- cism
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ – + 0 +
$TE_{a}(d)$ 0 0 0 0
[…]
A list of the normative effects for each form of fairness.
Table 3 lists the normative effects of each form of discrimination. By definition, the total effect must not be equal to zero for it to be considered discrimination: $TE_{a}(d)\neq 0$.
Table 3. Normative effects of discrimination
Effect “No” Reverse Covert Overt
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ + 0 0
$TE_{a}(d)$ + +
[…]
A list of the normative effects for each form of discrimination.
Notice that each form of fairness is associated with a different form of discrimination, leading to four fairness and discrimination pairs. That is, what might be considered fair for one group of people may be considered biased and discriminatory for another group of people. Fairness through supremacism (Figure 2a) and overt discrimination (Figure 2b) are both associated with the protected attribute $A$ having a negative direct effect on the decision $D$ ($A\rightarrow D$). Fairness through unawareness (Figure 2c) and covert discrimination (Figure 2d) are both associated with the protected attribute $A$ having no direct effect on the decision $D$. Fairness through affirmative action (Figure 2e) and reverse discrimination (Figure 2f) are both associated with the protected attribute $A$ having a positive direct effect on the decision $D$ ($A\rightarrow D$). Finally, fairness through lottery (Figure 2g) and “no” discrimination (Figure 2h) are both associated with the protected attribute $A$ having no direct effect on the decision $D$ and a positive indirect effect on the decision $D$ ($A\rightarrow W\rightarrow X\not\rightarrow D$).
5.1. Mapping fairness to politics
Table 4 shows how each form of fairness can be mapped to different positions on the political spectrum. Each position is associated with its own fairness criterion and fairness properties.
Table 4. Mapping fairness type to the political spectrum
Political worldview Far- left Left- wing Right- wing Far- right
Type Lottery Affirmative action Unaware- ness Suprem- acism
Fairness criterion $D\perp\!\!\!\perp A,W$ $D\perp\!\!\!\perp A$ $D\perp\!\!\!\perp A|W$ $D\>\neg\!\!\perp\!\!\!\perp A$
Individual fairness Yes No Yes No
Group fairness Yes Yes No No
Pearl’s law First Second
[…]
A table describing different forms of fairness with points on the political spectrum.
Right-wing notions of fairness rely on fairness through unawareness, which is associated with the criterion $D\perp\!\!\!\perp A|W$. That is, the decision $D$ is conditionally independent of the protected attribute $A$ given some mediator $W$. For example, a college admissions decision is independent of race given an applicant’s merit. This leads to individual fairness, where similar individuals are treated similarly regardless of protected group membership. This form of fairness relies on Pearl’s second fundamental law of causal inference: conditional independence.
Left-wing notions of fairness rely on affirmative action, which is associated with the criterion $D\perp\!\!\!\perp A$. That is, the decision $D$ is independent of the protected attribute $A$, which is achieved by using the protected attribute $A$ to correct for biases in the data $X$ that is used in the decision $D$. We imagine what the applicant’s outcome would have been had they been a member of a counterfactual class. We then perform an intervention on the decision to achieve the same outcome. For example, a college admissions decision would affirmatively use race to correct for deficiencies in the applicant’s merit due to structural discrimination. This leads to group fairness, where different groups receive more similar aggregate decision rates. This form of fairness relies on Pearl’s first fundamental law of causal inference: counterfactuals and interventions.
We now examine the two extreme forms of fairness. Far-right notions of fairness rely on fairness through supremacism, which is associated with the criterion $D\>\neg\!\!\perp\!\!\!\perp A$. In this case, individual and group notions of fairness are both violated. Far-left notions of fairness rely on fairness through lottery, which is associated with the criterion $D\perp\!\!\!\perp A,W$. In this case, individual and group notions of fairness are both satisfied.
5.2. Mapping discrimination to the law
Table 5 shows how each form of discrimination can be mapped to antidiscrimination law. There are two key legal concepts in antidiscrimination law. In the US, these are called disparate treatment and disparate impact. In the EU, these are called direct discrimination and indirect discrimination. Disparate treatment or direct discrimination occurs when an individual or group is treated differently based on a protected characteristic like race or gender. Disparate impact or indirect discrimination occurs when a seemingly neutral policy has a disproportionately adverse impact on a protected group.
Table 5. Mapping discrimination type to the law
US Type Disparate treatment Disparate impact
EU Type Direct discrimination Indirect discrimination
Overt X
Covert X
Reverse X
“No” X
[…]
A table describing different forms of discrimination with legal definitions of discrimination.
When a protected attribute $A$ has a direct effect on a decision $D$ ($A\rightarrow D$), we might consider this as a form of disparate treatment or direct discrimination. This would map to overt discrimination (Figure 2b) and reverse discrimination (Figure 2f). To avoid disparate treatment or direct discrimination, the protected attribute $A$ should have no direct effect on the decision $D$.
Related to disparate treatment is the concept of redlining, where a variable that is used in a decision $D$ acts as a proxy for a protected attribute $A$. For example, lenders have historically used zip code as a proxy for race as a basis for denying loans. We avoid redlining problems in this paper by only using data $X$ that is a direct measure of the mediator $W$. But a deeper discussion for how to prevent redlining using CBNs is discussed by Lam(Lam, 2024). He shows that certain types of alternative data can cause redlining because the protected attribute $A$ may act as a hidden confounder, leading to that data having a spurious effect on a decision $D$.
When a protected attribute $A$ has an indirect effect on a decision $D$, we might consider this as a form of disparate impact or indirect discrimination. This would map to covert discrimination (Figure 2d) and “no” discrimination (Figure 2h). Even thought the protected attribute $A$ has no direct effect on a decision $D$ (making the policy neutral), the inclusion of biased data or exclusion of unbiased data could cause the protected attribute $A$ to have an indirect effect on the decision $D$ that adversely affects a protected group.
For disparate impact or indirect discrimination, there is a three-step test. The first step is to establish that a neutral policy has a disproportionate impact on a protected group. This would involve testing whether there is a disparity in decisions $D$ across the protected class $A$. The second step is to establish that the business necessity of the policy. This would involve arguing that the data $X$ being used has some causal relationship to the mediator $W$. The final step is to identify a less discriminatory alternative model, which would involve performing a model search.
CBNs provide an analytical framework to reason about whether a model is causing illegal discrimination. By being able to explicitly model the different forms of discrimination, CBNs provide a common language that businesses and regulators can use to ensure that machine learning models are compliant with laws and regulations.
6. Modeling fairness as a causal loop diagram
One of the key limitations of causal Bayesian networks (CBNs) is that as directed acyclic graphs, they cannot be used for modeling feedback loops. But we need those feedback loops to understand the complex, dynamic effects of different fairness policies over time. So we turn our attention towards system dynamics, where we can leverage causal loop diagrams (CLDs) instead.
To model feedback loops within a CLD, we need to label each path as having a positive (+) or negative (–) direction. If the path is positive, then changes in one variable leads to changes in another variable in the same direction. For example, an increase in one variable would lead to an increase in the other variable. Alternatively, a decrease in one variable would lead to a decrease in the other variable. If the path is negative, then changes in one variable leads to changes in another variable in the opposite direction. For example, an increase in one variable would lead to a decrease in the other variable. Alternatively, a decrease in one variable would lead to an increase in the other variable.
CLDs are represented using a combination of reinforcing loops or balancing loops. Reinforcing loops compound changes over time, creating either virtuous cycles that cause growth or vicious cycles that cause decay. Balancing loops counteract changes over time, keeping a system within a fixed state that prevents growth or decay. Reinforcing loops have an even number of negative paths, while balancing loops have an odd number of negative paths.
We can combine reinforcing and balancing loops to create system archetypes. These represent common recurring patterns that explain counterintuitive behavior in complex systems. This concept was first coined by Senge(Senge, 1990) with additional research done by Meadows(Meadows, 2008) and Sterman(Sterman, 2000). Three system archetypes seem particularly helpful for understanding the fairness problem, which we have adapted from Kim(Kim, 1992).
The first archetype that we will examine is the “success to the successful” archetype, which is shown in Figure 3. This archetype consists of two reinforcing loops, R1 and R2. Let’s say that we have two equally capable groups that are competing for the same resources: groups A and B. If group A was initially given more resources, then that group will become more successful. That success provides justification to allocate even more resources to group A. This is shown by reinforcing loop R1. However, that could mean fewer resources are allocated to group B, which causes that group to become less successful. This results in fewer resources being allocated to group B. This is shown by reinforcing loop R2. Over the long run, R1 could create a virtuous cycle of growth for group A, while R2 could create a vicious cycle of decay for group B.
Allocation to Ainstead of BResourcesto AResourcesto BSuccessof ASuccessof BR1R2++++
Figure 3. Success to the successful archetype
[…]
Modeling fairness using the success to the successful archetype.
The next archetype that we will examine is the “limits to success” archetype, which is shown in Figure 4. This archetype consists of a reinforcing loop R1 and a balancing loop B1. Let’s say that a system’s performance continues to increase as a result of past efforts. This is shown by the reinforcing loop R1. But as performance increases, it begins to plateau as it reaches some limiting action. The limiting action is governed by some constraint that prevents further performance improvement. This is shown by the balancing loop B1. Even as efforts increase, the limiting action acts as a ceiling on performance.
[…]
Figure 4. Limits to success archetype
[…]
Modeling fairness using the limits to success archetype.
The final system archetype that we will examine is the “shifting the burden / addiction” archetype, which is shown in Figure 5. Let’s say that there is a problem symptom and two choices for a solution: an internal solution and an external intervention. The internal solution takes a long time to correct for the problem symptom (the double hash marks —— between the internal solution and the problem symptom represent a time delay), meaning that the problem symptom would continue to persist in the short term. On the other hand, the external intervention might alleviate the problem symptom immediately. But an unintended side effect of this solution is that it could create a dependence on the external intervention, which diverts attention away from an internal solution. This means that the problem symptom would actually persist over the long term, offsetting the benefit of the external intervention. This is called the shifting the burden archetype because the burden of solving the problem symptom is shifted from one party to another. It is also called an addiction archetype because people who rely on an external intervention may become addicted to that solution over the long term.
ProblemsymptomInternalsolutionExternalinterventionDependenceon externalinterventionB1B2R1$+$——
Figure 5. Shifting the burden / addiction archetype
[…]
Modeling fairness using the success to the successful archetype.
In Figure 6, we combine these three archetypes into a single CLD to model the algorithmic fairness problem. This diagram is essentially an expanded version of the CLD in Figure 1c. We begin with the “success to the successful” archetype. A machine learning model is embedded inside decision $D$, which makes allocation decisions based on decision $X$ and has a tendency to favor one group over another. Those decisions lead to resources being assigned to each group, which eventually determines that group’s success and is measured as an outcome $Y$. Past success $Y$ becomes future data $X$, which is then fed back into the machine learning model to drive a future allocation decision $D$. This is represented using reinforcing loops R1 and R2. Since the allocation decisions $D$ are initially only driven by data $X$, success to the successful is equivalent to fairness through unawareness (Figure 2c).
Allocation to Ainstead of B ($D$)Resourcesto AResourcesto BSuccessof A ($Y_{A}$)Successof B ($Y_{B}$)Data fromA ($X_{A}$)Data fromB ($X_{B}$)R1R2Limitingaction on ALimitingaction on BB1B2Constrainton AConstrainton BExternalinterventionby A ($A_{A}$)Externalinterventionby B ($A_{B}$)Dependenceon externalinterventionby ADependenceon externalinterventionby BMLModelB3B4R3R4+++++++++++————
Figure 6. Causal loop diagram representation of algorithmic fairness
[…]
Modeling fairness using the success to the successful, limits to success, and the shifting the burden / addiction archetypes.
We next add in the “limits to success” archetype. As resources to group A or group B increases, it may hit up against a limiting action that is governed by come constraint. This then reduces the success of that group as measured by outcome $Y$. As the group’s success becomes limited, this is reflected in the data $X$ for that group. This leads to a reduction in the allocation decision $D$ of resources to the previously successful group. This is represented by balancing loops B1 and B2. Whereas success to the successful causes a previously successful group to become more successful, limits to success may give a less successful group a chance to catch up.
Finally, we add in the “shifting the burden / addiction” archetype. If one group is allocated less resources than another group, then it could breed resentment by the disadvantaged group. That group could then demand that the protected attribute $A$ be used to externally intervene on the allocation decision $D$. This external intervention is equivalent to fairness through affirmative action (Figure 2e). While it might reduce allocation disparities in the short term, it may actually create a dependence on the external intervention over the long term. This is because if allocation decisions $D$ are not based purely on data $X$, which is a measure of mediator $W$ (e.g. merit, qualifications), then the assignment of resources to the disadvantaged group may not lead to their success $Y$. For example, approving a loan to a disadvantaged applicant even though the applicant is not creditworthy could lead to a worse outcome (default) than if the loan was denied in the first place. That decreased success in outcome $Y$ for the less successful group could trigger a demand for an even stronger external intervention $A$ on the decision $D$, reinforcing this problem. In the meanwhile, attention may be taken away from the internal solution of optimizing for the mediator $W$ and the data $X$ so that the disadvantaged group would perform better under fairness through unawareness. This may especially be the case if the disadvantaged group believes that the mediator $W$ is biased and therefore does not focus on optimizing that metric.
Whereas CBNs are very effective at modeling the normative effects of fairness and discrimination, CLDs are very effective at modeling the positive effects. Recall from Hume’s law that positive effects model what is, while normative effects model what ought to be (i.e. the is–ought problem).
Table 6 lists the positive effects of the different forms of fairness and discrimination. As mentioned before, each form of fairness is associated with a form of discrimination, leading to the same direct effect $DE_{a}(d)$. Whereas normative effects differ based on our prior beliefs about where bias exists in a model, positive effects do not take into account our prior beliefs of bias. However our analysis of positive effects depends on whether we analyze disparities across groups from a short-term or a long-term perspective. Notice that for fairness through supremacism or overt discrimination, since a disadvantaged group is being overtly disadvantaged, there is a doubly negative total effect ($TE_{a}(d)\ll 0$).
Table 6. Positive effects of fairness and discrimination
Fairness Lottery Affirmative action Unaware- ness Suprema- cism
Discrim- ination “No” Reverse Covert Overt
Short-term
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ 0
$TE_{a}(d)$ 0 0 – –
Long-term
$DE_{a}(d)$ 0 + 0
$IE_{a}(d)$ 0 – – – +
$TE_{a}(d)$ 0 0 – –
[…]
A list of the positive effects for each form of discrimination.
During the short-term, we only examine the “success to the successful” archetype as shown in Figure 6 through feedback loops R1 and R2. A policy of fairness through unawareness would lead to a disadvantaged group receiving worse decisions than an advantaged group ($TE_{a}(d)<0$). This is because fairness through unawareness would replicate and reinforce the past success of the advantaged group over the disadvantaged group ($IE_{a}(d)<0$). But an affirmative action policy could help to equalize the decision rates across both groups. Therefore, the total effect from fairness through unawareness is negative ($TE_{a}(d)<0$) and the total effect from affirmative action is zero ($TE_{a}(d)=0$).
During the long-term, we add in the effects of the “limits to success” archetype as shown in Figure 6 through feedback loops B1 and B2. This archetype puts a constraint on how successful the advantaged group can become, reducing the disparities between the advantaged group and the disadvantaged group ($IE_{a}(d)>0$). We also add in the effects of the “shifting the burden / addiction” archetype as shown in Figure 6 through feedback loops R3, B3, R4, and B4. This archetype can create a dependency by the disadvantaged group on the external intervention and harm their long-term success ($IE_{a}(d)<0$). Therefore, the total effect from fairness through unawareness is zero ($TE_{a}(d)=0$) and the total effect from affirmative action is negative ($TE_{a}(d)<0$). That is, the total effects between the short-term and long-term versions of these fairness policies may flip direction.
6.1. Case study: Racial disparities in the US
We can use the CBNs from Figure 2 and the CLD from Figure 6 to model the complex political debate over racial disparities in the US. According to the US Census Bureau, median household income in 2021 was \$100,573 for Asians, \$74,932 for Whites, \$57.671 for Hispanics, and \$48,297 for Blacks(Survey, 2021). Meanwhile, median household wealth in 2021 was \$320,900 for Asians, \$250,400 for Whites, \$48,700 for Hispanics, and \$27,100 for Blacks(Kochnar and Moslimani, 2023). We would like to understand why Whites and Asians have higher median household income and wealth than Blacks and Hispanics.
When it comes to understanding the root causes of racial disparities, we can take into account some of the original sources of bias in the data. We know that the US has historically had unjust policies that overtly discriminated against non-Whites. These included racist policies like slavery, Jim Crow segregation, and the Chinese Exclusion Act. For convenience, we replicate the CBN for overt discrimination in Figure 7.
[…]
Figure 7. Overt discrimination
[…]
A CBN with where the protected attribute $A$ has a negative direct effect on the decision $D$.
Overt discrimination causes bias in both the decision $D$ and the outcome $Y$. The outcome $Y$ becomes future data $X$ due to a feedback loop. Figure 8 shows what this feedback loop likes within a CBN. We temporarily violate the directed acyclic graph constraint of CBNs and create a path $Y\rightarrow W$, causing the mediator $W$ to become biased.
[…]
Figure 8. Feedback from overt discrimination
[…]
A CBN with a feedback loop from the outcome $Y$ to the mediator $W$.
The bias in mediator $W$ gets replicated in the data $X$, which then causes bias in future decisions $D$ and outcomes $Y$. Even though the US passed the Civil Rights Act of 1964 banning overt discrimination, this didn’t reverse the initial bias caused by historically unjust policies against non-Whites. Even a racially neutral policy could still form a basis for covert discrimination as shown in Figure 9.
[…]
Figure 9. Covert discrimination
[…]
A CBN with where the protected attribute $A$ has a negative indirect effect on the decision $D$.
This creates a dilemma for policymakers: Should the US institute fairness through unawareness, even if such a policy could replicate past racial disparities (i.e. the “success to the successful” archetype)? Or should the US institute affirmative action to try to correct for past injustices? Let’s say that US policymakers (particularly Democrats) succeed in instituting affirmative action as shown in Figure 10.
[…]
Figure 10. Fairness through affirmative action
[…]
A CBN with where the protected attribute $A$ has a positive direct effect on the decision $D$.
While affirmative action might help correct for biases in the decision $D$, notice in Figure 10 that it does not necessarily correct for biases in the outcome $Y$. The protected attribute $A$ cannot have a direct effect on the outcome $Y$. It can only directly affect the decision $D$. Correcting for biases in the decision $D$ cannot fully offset biases in the mediator $W$, which means that the outcome $Y$ remains biased. The path $D\rightarrow Y$ only has a weak effect on the outcome $Y$. On the other hand, the path $W\rightarrow Y$ has a strong effect on the outcome $Y$. This is shown in Figure 11.
[…]
Figure 11. Strong and weak effects
[…]
A CBN where the outcome $Y$ receives a strong direct effect from mediator $W$ and a weak direct effect from the decision $D$.
For example, an affirmative action policy in college admissions might help more Blacks and Hispanics to be accepted ($A\rightarrow D$). But because they were not accepted purely based on their merit (mediator $W$), they might be less likely to graduate (outcome $Y$). This is because they might be put into a class where their White and Asian counterparts would consistently outperform them. Another example could be an affirmative action policy for credit decisions. This might help more Blacks and Hispanics to receive more loans ($A\rightarrow D$). But since they were not accepted purely based on their creditworthiness (mediator $W$), they might be more likely to default (outcome $Y$). This is because affirmative action does not directly improve one’s merit or creditworthiness. Even worse, affirmative action could have the opposite effect on the mediator $W$.
Due to the “shifting the burden / addiction” archetype, an affirmative action policy may reinforce a dependency on this external intervention and take away attention from an internal solution. Figure 12 shows that the internal solution would be to focus on reducing disparities in the mediator $W$. But instead, affirmative action shifts attention to reducing disparities in the decision $D$.
[…]
Figure 12. Internal solution and external intervention
[…]
A CBN where disparities in outcomes are either closed by modifying the mediator $W$ or the decision $D$.
This creates a major problem. Affirmative action may create a false illusion that you can have a system of double standards and not pay the price. While it could enable less qualified Blacks and Hispanics to get access to the some of the same resources that would have went to more qualified Whites and Asians, it doesn’t necessarily translate into better overall outcomes. In college admissions for example, one could argue that a Black or Hispanic college student struggling at an elite university might have been better off thriving at a flagship state university instead. In employment, a Black or Hispanic employee might have been better off being placed in a team where they were at the same level as their White and Asian peers versus being placed in a higher performing team where they ended up dragging the rest of the team down. This is the main argument that has been advanced by Sowell, Mac Donald, and other conservative scholars(Sowell, 2004, 2019; Donald, 2023).
Placing Blacks and Hispanics into higher level positions than is warranted by their merits and qualifications could also exact a harmful psychological toll. This could reinforce their views that they are victims of an unjust system, leading to a sense of learned helplessness and a loss of agency. This could result in reduced productivity, which is the primary determinant of income and wealth in the capitalist economic system of the US. Rather than helping Blacks and Hispanics become more competitive against Whites and Asians, affirmative action may have instead made them less competitive.
Asians in the US have not historically benefited from affirmative action policies. Asians have instead relied on fairness through unawareness, as shown in Figure 13. In some cases like college admissions, they may even have to outperform their White counterparts to receive the same admission decisions(Grossman et al., 2023). Asians as a group have tended to focus their efforts on the internal solution of competing based on their merits and qualifications. By working against Asians, affirmative action programs may have actually caused Asians to work even harder than other racial groups. By focusing on developing their human capital, not only have Asians closed the gap in income and wealth with Whites over the long-term, they have actually surpassed them. Over time, this has also reduced perceived biases in the data between Asians and Whites.
[…]
Figure 13. Fairness through unawareness
[…]
A CBN with where the protected attribute $A$ has no effect on the decision $D$.
This is the reason why some US policymakers (particularly Republicans) have been arguing in favor of fairness through unawareness. While acknowledging that historically unjust policies like slavery and Jim Crow have created the initial disparities between Whites and non-Whites, they also recognize that external government interventions like affirmative action may cause more harm than good. They would argue that it is not an accident that the racial group that has benefited the most from affirmative action (i.e. Blacks) ultimately received the worst outcomes. Meanwhile, the racial group that has benefited the least from affirmative action (i.e. Asians) ultimately received the best outcomes.
That being said, this doesn’t mean that affirmative action would not have any value. If it can be used to identify and correct for false positives, then affirmative action might be a helpful way to marginally reduce disparities across groups. The key here is to use the protected attribute $A$ to only correct for biases in the data $X$ that may overpredict the probability of a bad outcome $Y$, as opposed to using it to directly achieve some type of diversity quota. We would need to ensure that an affirmative action policy that reduces disparities in the decision $D$ does not result in an increase in the disparities in the outcome $Y$.
Affirmative action is a low-leverage solution that should only be used to modestly decrease racial disparities. It is not a fundamental, high-leverage solution for reducing racial disparities over the long-term. As long as there are racial disparities in the mediator $W$, then there will be racial disparities in the outcome $Y$. This would mean that racial disparities in household income and wealth in the US would continue to persist. The focus should instead be on closing disparities in the mediator $W$. For example, Democrats could call for increased government funding of education and training programs. On the other hand, Republicans could call for the banning of racial preferences in high-stakes decisions.
Hopefully, these models provide a basis for arguing that while historical policies like slavery and Jim Crow generated the initial racial disparities between Whites and Blacks, contemporary policies like affirmative action and DEI may have reinforced those disparities. Affirmative action should only be used as a temporary solution for marginally closing the racial equity gap and if it is used, then there should be a long-term plan to eventually switch back to fairness through unawareness. If an addictive dependency on affirmative action develops that diverts attention away from internal solutions for closing the merit and qualifications gap, then the affirmative action policy should be ended. Systems thinking shows us that AI researchers should not be so quick to dismiss fairness through unawareness as a fundamental solution for closing the racial equity gap.
7. Modeling fairness as a causal hierarchy
Systems thinking allows us to combine three previously independent analytical techniques: machine learning, causal inference, and system dynamics. Doing so allows us to capture the emergent behavior from three different aspects of the fairness problem. We can think of causal Bayesian networks as a higher level representation of supervised machine learning models, which emerges when we “think outside of the black box.” We can think of causal loop diagrams as a higher level representation of causal Bayesian networks, which emerges when we violate the directed acyclic graph constraint and create a feedback loop.
However, this suggests the need for a new causal hierarchy to model the algorithmic fairness problem. The left side of Figure 14 shows the causal hierarchy proposed by Pearl, which is not optimized for supervised machine learning and does not incorporate system dynamics. The right side of Figure 14 is a proposed new causal hierarchy that more explicitly models the emergent behavior between supervised machine learning, causal inference, and system dynamics.
Layer 1. Association
Activity:
Seeing, observing
Question: What if I see…? Example: What does a symptom tell me about a disease?
Layer 2. Intervention
Activity:
Doing, intervening
Question: What if I do…? How? Example: If I take aspirin, will my headache be cured?
Layer 3. Counterfactuals
Activity:
Imagining, retrospection, understanding
Question: What if I had done…? Why? Example: Was it the aspirin that stopped my headache?
Layer 1. Prediction
Field:
Supervised machine learning
Representation: Block diagrams Type: Association
Layer 2. Control
Field:
Causal inference (Judea Pearl)
Representation: Causal Bayesian networks Type: Acyclic causation
Layer 3. Simulation
Field:
System dynamics (Jay Forrester)
Representation: Causal loop diagrams Type: Cyclic causation
[…]
Figure 14. Causal hierarchy representation of algorithmic fairness
[…]
Adding a system dynamics layer to Pearl’s causal hierarchy.
We can directly map layer 1 of Pearl’s causal hierarchy (association) to layer 1 of this new causal hierarchy (association), but we explicitly define this layer to be used for supervised machine learning. We can then combine layers 2 and 3 of Pearl’s causal hierarchy (intervention and counterfactuals) into layer 2 of this new causal hierarchy (acyclic causation). Finally, we add a new layer 3 to the causal hierarchy for system dynamics (cyclic causation).
Each of the layers in this new causal hierarchy have unique properties and behaviors that make them optimized for different applications. The first layer is associated with supervised machine learning and is optimized for prediction. The second layer is associated with causal inference and is optimized for control. The third layer is associated with system dynamics and is optimized for simulation. These three new causal layers – prediction, control, and simulation – make up an expanded causal hierarchy that is consistent with systems thinking principles. It may provide us with a new sociotechnical approach for evaluating how AI systems interact with society.
8. Modeling fairness as a systems map
In this final section, we will combine all of the ideas from the previous sections into a very high level representation of the algorithmic fairness problem called a systems map. This is shown in Figure 15. On the left side of the systems map, we use a connection circle to show the different fields in the social sciences that are related to fairness. These ideas were described in greater detail in section 3. On the right side of the systems map, we use a causal hierarchy to show the different technical aspects of the fairness problem. These ideas were described in greater detail in section 7.
Social Technical Psychology External Internal
attribution
attribution
Sociology Structure Agency Economics Socialism Capitalsm Anthropology Social Cultural Politics Left-wing Right-wing Law Disparate Disparate
impact
treatment
System Dynamics Success Shifting the
to the
burden /
successful
Addiction
Causal Inference First law: Second law:
Counterfactuals
Conditional
and
independence
interventions
Machine Learning Fairness Fairness
through
through
affirmative
unawareness
action
CausalBridgeRefer to caption
Figure 15. System map representation of algorithmic fairness
[…]
System map showing the causal links between the social sciences and computer science.
As mentioned before, algorithmic fairness is a sociotechnical problem. We need to find a way to bridge the social and technical aspects of the fairness problem. In each of the social science fields, there exists a dichotomy that represents different viewpoints for understanding the fairness problem. This allows us to understand the fundamentally ideological and political debate over fairness, which centers around how we attribute the causes for an outcome. The systems map shows a causal bridge that links this dualism in the social sciences to Pearl’s two fundamental laws of causal inference, which was discussed in section 5.1. Hopefully this systems map provides a very compact and high level representation for how the different fields of fairness relate to each another.
9. Conclusion
Systems thinking provides us with a novel way to model the algorithmic fairness problem by “thinking outside of the black box.” While a large amount of AI research has been focused on the hidden layers that are “inside” a black box ML model (e.g. neural networks), this paper has been focused on the transparent layers that are “outside” a black box ML model. These outer layers are where we can encode our knowledge and assumptions about where we might believe that bias occurs in the data generating process. We showed that the first outer layer can be represented using causal Bayesian networks and the second outer layer can be represented using causal loop diagrams. This gives us an elegant way of modeling the emergent behavior between machine learning, causal inference, and system dynamics.
This approach allow us to develop a more unified and holistic view of the fairness and machine learning problem, providing the essential connections between computer science and the law. This is the key to building AI systems that are aligned with society’s democratic values and the policymakers’ political agendas. It will enable businesses and regulators to communicate with each other using a common causal language, one that can be used by both data scientists and lawyers. These models may even aid in AI deregulation by identifying mathematical flaws in the existing laws and regulations that are caused by paradoxes. This will help businesses to comply with antidiscrimination laws and regulations, reducing business uncertainty and expensive litigation while providing a much needed layer of transparency and trust for consumers of AI products and services.
Whereas much of the fairness and machine learning literature has been focused on how fairness through unawareness perpetuates racial inequalities, this paper demonstrates how affirmative action and DEI may also perpetuate racial inequalities. It shows us that racially aware policies are not a silver bullet in closing the racial income and wealth gap. We can use systems thinking to move away from shallow and ideological notions of fairness and towards deeper and holistic notions of fairness that reflect how the real world actually works. We argue that racial disparities in the US are not purely due to historical, structural, and systemic factors. Personal responsibility plays as important of a role (if not more). Therefore, instead of just focusing on eliminating structural and systemic barriers to racial equity, we need to put as much attention towards creating systems that motivate individuals to make good decisions while holding them accountable for making bad decisions.
Systems thinking supplies the missing mathematical models that are needed to build responsible AI systems. To solve fairness, we need to bring together both the social and technical aspects of what is ultimately a complex systems problem. On the social side, it requires a diversity of perspectives that is inclusive of people from both sides of the political spectrum. On the technical side, it requires unifying analytical techniques that were previously thought of as being independent. By unifying all of these concepts together, we hope to build AI policies and systems that could lead to actual, real-world solutions for closing disparities across society.